OpenStudy (anonymous):

Derivatives. y= 1/ (1-x^2)

OpenStudy (anonymous):

\[(\frac{1}{f})'=\frac{-f'}{f^2}\] in this case you get \[\frac{2x}{(1-x^2)^2}\]